OpenStudy (anonymous):

A horse race has 13 entries. Assuming that there are no ties, what is the probability that the four horses owned by one person finish first, second, third and fourth?

OpenStudy (anonymous):

probability of that is = 9! / 13! = 1/ (13 * 12 * 11 * 10) = 1/ 17160